# mathx”00 mathx”08 mathx”10 mathx”18 mathx”20 mathx”28 mathx”30 mathx”38 mathx”40 mathx”44

### v1

@@ 1.3 The model
We study the sine-Gordon measures. Let $\beta\in[4\pi,8\pi)$ and $\zeta$ be a bounded Borel function on the unit torus $\mathbb{T}^{2}$. [^2] The restriction on $\beta$ away from 0 is not necessary, we have made it to avoid constants depending on $\beta^{-1}$ which seems to only be a distraction. The interested reader can easily modify the results herein to cover this case, which can anyway be treated by other methods giving stronger results. Furthermore pick $\varepsilon>0$ and then let $\gamma[C_{\varepsilon}]$ be the law of the mollified unit mass GFF $\chi_{\varepsilon}*\Phi$ (where $\chi_{\varepsilon}$ is a standard mollifier). [^3] We put in a mass to keep things as simple as possible, primarily to avoid issues with the zero mode which arise in the massless model. Setting it equal to $1$ is convenient because it equals the width of the torus. Of course, this is not a “generic” choice, but the additional complications in case the mass is not fixed to the torus size did not seem illuminating enough to warrant investigation in this note. Put
@@
$\mu_{\mathrm{sG}}$ does make sense because $\mathrm{C}_{\varepsilon}$ is a UV-mollified GFF covariance with cutoff $\varepsilon$ so that the Gaussian measure $\gamma[\mathrm{C}_{\varepsilon}]$ is $\mathcal{C}^{\infty}$-valued. If $\mathrm{T}$ is a symmetry of the torus,
@@
At this point we fix some constants. First, $\sigma\coloneqq 1-\frac{\beta}{8\pi}\in(0,\frac{1}{2}]$ is the distance to criticality. $\delta$ is a small positive number which could be fixed at, say, $\delta=\frac{1}{100}$ throughout, $\ell\gtrsim_{\sigma^{-1},\delta^{-1}}1$ is an integer, $L=\ell^{\mathfrak{m}}$ where $\mathfrak{m}=\ell^{2}$, $A\gtrsim_{L}1$, and $\zeta_{*}\lesssim_{A^{-1}}1$. We also introduce $\mathfrak{c}^{(1)},\mathfrak{c}^{(2)},\mathfrak{c}^{(3)},\mathfrak{c}^{(4)}$ with $\mathfrak{c}^{(2)}\lesssim_{1}1\lesssim_{1}\mathfrak{c}^{(1)}\wedge\mathfrak{c}^{(3)}$, $\mathfrak{c}^{(4)}\gtrsim_{1}\mathfrak{c}^{(3)}$, and $\mathfrak{h}_{*}\gtrsim_{1}1$, $\kappa_{*}=\mathfrak{c}^{(4)}\log L$. $\zeta_{\bullet}$ is a completely arbitrary positive number and will represent the upper bound on the size of the coupling $\zeta$ as measured in $\mathrm{L}^{\infty}$. $j_{\bullet}$ is a negative integer defined as
@@ 2.1 Progressive integration
Now we define
@@ Appendix A Scale decomposition
A.1 Continuous decomposition
@@
${\mathrm{F}}_{\hskip-1.0ptt}(x;\hskip-0.5ptm)=t\sum_{p\in 2\pi\mathbb{Z}}\varphi\big{(}t\sqrt{m^{2}+p^{2}}\big{)}e^{ipx}$ (80)
@@
Let $\chi$ be the standard smooth bump function on $\mathbb{R}^{2}$ with support radius $1$. For $\varepsilon>0$, put $\chi_{\varepsilon}(x)=\varepsilon^{-2}\chi(\varepsilon^{-1}x)$. After mollifying and periodizing, we can decompose
@@
${\mathrm{F}}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)$ $=(\chi_{\varepsilon}*{\mathrm{F}}_{\hskip-1.0ptt})(x)=t\hskip 1.0pt\sum_{p\in 2\pi\mathbb{T}^{2}}\hat{\chi}(\varepsilon p)\varphi\big{(}t\sqrt{m^{2}+p^{2}}\big{)}e^{ipx}.$ (82)
@@
The following estimate is immediate for any multiindex $\alpha$ (it holds for $\varepsilon=0$ as well):
$|\partial^{\alpha}_{x}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)|\lesssim_{|\alpha|}t^{-|\alpha|}e^{-\sqrt{tm}}.$ (84)
@@
$|\partial^{\alpha}_{x}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)|\lesssim_{|\alpha|}(\varepsilon^{-1}t)^{2}\varepsilon^{-|\alpha|}.$ (85)
@@
A.2 Discrete decomposition & scaling
@@
$|\partial^{\alpha}\Gamma_{\hskip-1.0pt\varepsilon;0}(x)|\lesssim_{|\alpha|}1.$ (94)
@@
$|\partial^{\alpha}(\Gamma_{\hskip-1.0pt\varepsilon;j}-\Gamma_{\hskip-1.0pt\dot{\varepsilon};j})|(x)$ $\lesssim_{|\alpha|}L^{-j}|\varepsilon-\dot{\varepsilon}|,$ (95)
@@
A.3 Finer scale decomposition
@@ C.1 Algebraic inequalities
where we have had to require $\mathfrak{c}^{(1)}\gtrsim_{1}\mathfrak{c}^{(2)}$. Moving on, by (135,140) and the triangle inequality,
@@
Now if $X$ is small and the conditions on constants as above, take (163) and combine it with (139,145):
@@ C.2 Fluctuation step bounds
$\mathbb{E}\big{[}\!\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\|\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s}\|^{2}_{\mathrm{L}^{2}(X)}\big{]}\leqslant\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\mathrm{Tr}\,\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})).$ (172)
@@
By assumption, $\mathrm{Rem}_{0}\mathbb{E}\big{[}k^{(0)}_{j}(X,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}=\mathrm{Rem}_{1}\mathbb{E}\big{[}k^{(0)}_{j}(X,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}$. From (185,186),
@@ Appendix E Laplace transforms
Let $\mathfrak{F}$ be the Fourier transform, which induces an isomorphism of the dual pairs $(\mathcal{C}^{\infty}(\mathbb{T}^{d}),\mathcal{D}^{\prime}(\mathbb{T}^{d}))$ and $(\mathcal{S}_{\mathrm{i}}^{\vphantom{a}}(\mathbb{Z}^{d};\mathbb{C}),\mathcal{S}_{\mathrm{i}}^{\prime}(\mathbb{Z}^{d};\mathbb{C}))$. The second couple are just function spaces on $\mathbb{Z}^{d}$ with the natural polynomial growth conditions, where the subscript indicates that the functions in question satisfy $f(-x)=\overline{f(x)}$. The duality pairing is explicitly given by
@@ Acknowledgments
I thank Roland Bauerschmidt, Jon Dimock, Massimiliano Gubinelli, Peter Paulovics, and Sarah-Jean Meyer for useful input. Special thanks go to David Brydges for his help and many enlightening conversations without which this paper might have appeared. The work was supported by Oxford’s Clarendon Fund.

### v2

@@ 1.3 The model
We study the sine-Gordon measures. Let $\beta\in[4\pi,8\pi)$ and $\zeta$ be a bounded Borel function on the unit torus $\mathbb{T}^{2}$. [^2] The restriction on $\beta$ away from 0 is not necessary, we have made it to avoid constants depending on $\beta^{-1}$ which seems to only be a distraction. The interested reader can easily modify the results herein to cover this case, which can anyway be treated by other methods giving stronger results. Furthermore pick $\varepsilon>0$ and then let $\gamma[C_{\varepsilon}]$ be the law of the mollified unit mass [^3] We put in a mass to keep things as simple as possible, primarily to avoid issues with the zero mode which arise in the massless model. Setting it equal to $1$ is convenient because it equals the width of the torus. Of course, this is not a “generic” choice, but the additional complications in case the mass is not fixed to the torus size did not seem illuminating enough to warrant investigation in this note. GFF $\chi_{\varepsilon}*\Phi$ ($\chi_{\varepsilon}$ is a not-too-smooth radially symmetric compactly supported mollifier with integral $1$, see Appendix A.1). Put
@@
$\mu_{\mathrm{sG}}$ does make sense because $\mathrm{C}_{\varepsilon}$ is a UV-mollified GFF covariance with cutoff $\varepsilon$ so that the Gaussian measure $\gamma[\mathrm{C}_{\varepsilon}]$ is at least $\mathcal{C}^{0}$-valued. If $\mathrm{T}$ is a symmetry of the torus,
@@
At this point we fix some constants. First, $\sigma\coloneqq 1-\frac{\beta}{8\pi}\in(0,\frac{1}{2}]$ is the distance to criticality. $\mathfrak{d}$ is an integer of order (say) $10$ and determines how many derivatives $\chi_{\varepsilon}$ has; equivalently, its decay in momentum space. $\delta$ is a small positive number which could be fixed at, say, $\delta=\frac{1}{100}$ throughout, $\ell\gtrsim_{\sigma^{-1},\delta^{-1}}1$ is an integer, $L=\ell^{\mathfrak{m}}$ where $\mathfrak{m}=\ell^{2}$, $A\gtrsim_{L}1$, and $\zeta_{*}\lesssim_{A^{-1}}1$. We also introduce $\mathfrak{c}^{(1)},\mathfrak{c}^{(2)},\mathfrak{c}^{(3)},\mathfrak{c}^{(4)}$ with $\mathfrak{c}^{(2)}\lesssim_{1}1\lesssim_{1}\mathfrak{c}^{(1)}\wedge\mathfrak{c}^{(3)}$, $\mathfrak{c}^{(4)}\gtrsim_{1}\mathfrak{c}^{(3)}$, and $\mathfrak{h}_{*}\gtrsim_{1}1$, $\kappa_{*}=\mathfrak{c}^{(4)}\log L$. $\zeta_{\bullet}$ is a completely arbitrary positive number and will represent the upper bound on the size of the coupling $\zeta$ as measured in $\mathrm{L}^{\infty}$. $j_{\bullet}$ is a negative integer defined as
@@ 2.1 Progressive integration
This is legitimate because $\mathrm{C}_{\varepsilon}$ has polynomial decay in momentum space, hence any smooth $J$ is in its Cameron-Martin space. Now we define
@@ Appendix A Scale decomposition
A.1 Mollifier
We take $\chi=c_{\mathfrak{d}}\mathsf{B}^{*2}\mathrm{C}^{*\mathfrak{d}}$ where $\mathsf{B}$ is a standard bump of support radius $1/4$ on $\mathbb{R}^{2}$ and $c_{\mathfrak{d}}$ is chosen so that $\chi$ has integral $1$. In Fourier space,
$\hat{\chi}(p)\propto_{1}\int_{\mathbb{R}^{2}}\frac{\hat{\mathsf{B}}(p-q)^{2}}{(1+|q|^{2})^{\mathfrak{d}}}\mathrm{d}q.$ (79)
Because $\hat{\mathsf{B}}$ is Schwartz, it is easy to see that this still behaves asymptotically like $(1+|p|^{2})^{\mathfrak{d}}$. For $0<\varepsilon<1$, put $\chi_{\varepsilon}(x)=\varepsilon^{-2}\chi(\varepsilon^{-1}x)$ (this also makes sense as a function on $\mathbb{T}^{2}$).
A.2 Continuous decomposition
@@
${\mathrm{F}}_{\hskip-1.0ptt}(x;\hskip-0.5ptm)=t\sum_{p\in 2\pi\mathbb{Z}^{2}}\varphi\big{(}t\sqrt{m^{2}+p^{2}}\big{)}e^{ipx}$ (81)
@@
After mollifying and periodizing, we can decompose
@@
${\mathrm{F}}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)$ $=(\chi_{\varepsilon}*{\mathrm{F}}_{\hskip-1.0ptt})(x)=t\hskip 1.0pt\sum_{p\in 2\pi\mathbb{Z}^{2}}\hat{\chi}(\varepsilon p)\varphi\big{(}t\sqrt{m^{2}+p^{2}}\big{)}e^{ipx}.$ (83)
@@
The following estimate is immediate for any multiindex $\alpha\leqslant 4$ and $\varepsilon\geqslant 0$:
$|\partial^{\alpha}_{x}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)|\lesssim_{1}t^{-|\alpha|}e^{-\sqrt{tm}}.$ (85)
@@
$|\partial^{\alpha}_{x}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)|\lesssim_{1}(\varepsilon^{-1}t)^{2}\varepsilon^{-|\alpha|}.$ (86)
@@
A.3 Discrete decomposition & scaling
@@
$|\partial^{\alpha}\Gamma_{\hskip-1.0pt\varepsilon;0}(x)|\lesssim_{1}1.$ (95)
@@
$|\partial^{\alpha}(\Gamma_{\hskip-1.0pt\varepsilon;j}-\Gamma_{\hskip-1.0pt\dot{\varepsilon};j})|(x)$ $\lesssim_{1}L^{-j}|\varepsilon-\dot{\varepsilon}|,$ (96)
@@
A.4 Finer scale decomposition
@@ C.1 Algebraic inequalities
where we have had to require $\mathfrak{c}^{(1)}\gtrsim_{1}\mathfrak{c}^{(2)}$. Moving on, by (136,141) and the triangle inequality,
@@
Now if $X$ is small and the conditions on constants as above, take (164) and combine it with (140,146):
@@ C.2 Fluctuation step bounds
$\mathbb{E}\big{[}\!\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\|\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s}\|^{2}_{\mathrm{L}^{2}(X)})\big{]}\leqslant\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\mathrm{Tr}\,\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})).$ (173)
@@
By assumption, $\mathrm{Rem}_{0}\mathbb{E}\big{[}k^{(0)}_{j}(X,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}=\mathrm{Rem}_{1}\mathbb{E}\big{[}k^{(0)}_{j}(X,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}$. From (186,187),
@@ Appendix E Laplace transforms
Let $\mathfrak{F}$ be the Fourier transform, which induces an isomorphism of the dual pairs $(\mathcal{C}^{\infty}(\mathbb{T}^{d}),\mathcal{D}^{\prime}(\mathbb{T}^{d}))$ and $(\mathcal{S}_{\mathrm{i}}^{\vphantom{a}}(\mathbb{Z}^{d};\mathbb{C}),\mathcal{S}_{\mathrm{i}}^{\prime}(\mathbb{Z}^{d};\mathbb{C}))$. The second couple are just function spaces on $\mathbb{Z}^{d}$ with the natural polynomial growth conditions, where the subscript indicates that the functions in question satisfy $f(-x)=\overline{\hskip-1.0ptf(x)\hskip-1.0pt}\,$. The duality pairing is explicitly given by
@@ Acknowledgments
I thank Roland Bauerschmidt, Jon Dimock, Massimiliano Gubinelli, Peter Paulovics, and Sarah-Jean Meyer for useful input. Special thanks go to David Brydges for his help and many enlightening conversations without which this paper might not have appeared. The work was supported by Oxford’s Clarendon Fund.